NarutoandSasuke:

can someone help with finding the area of a triangle 2x-2 height x+8 base (multiplying polynomials)

SkyVoltage43:

@HuskyNation

SkyVoltage43:

give me assistance

HuskyNation:

hmmm never understood this content when i did it sorry ;-;

SkyVoltage43:

darn

MARC:

\(Area~of\triangle=\frac{1}{2}\times~base\times~height\)

MARC:

\(=\frac{1}{2}\times(x+8)\times(2x-2)\)

MARC:

\(=\frac{1}{2}\times(2x^2-2x+16x-16)\)

MARC:

\(=\frac{1}{2}\times(2x^2+14x-16)\)

MARC:

\(=\frac{1}{2}\times2(x^2+7x-8)\)

MARC:

\(=?\) solve it @NarutoandSasuke

SkyVoltage43:

lol marc my buddy already found it out i think

MARC:

oh I c... x'D

NarutoandSasuke:

Thanks guys i got it.

MARC:

np

SkyVoltage43:

@NarutoandSasuke if you ever need math help just go to marc he is ggod